Suppose that in a certain restaurant the number of hamburgers sold varies inversely as the price. If 70 burgers are sold at $1.50 each, how many burgers would be sold at $1.25 each?
Answer by fractalier(6550)

We write this inverse relationship as
y = k/x
or in this case
n = k/p
now plug in to find k
70 = k/1.50
k = 105
now rewrite it using the found k
n = 105/p
and plug in the new data to find n
n = 105/1.25
n = 84